OpenStudy (anonymous):

I have to write a linear factorization of this: f(x) = x4 + 64x2

OpenStudy (loser66):

hey, this problem is not hard for you for sure

OpenStudy (anonymous):

These are my choices: A) f(x) = x^2(8x + i)(8x - i) B) f(x) = x^2(x + 8i)^2 C) f(x) = x^2(x + 8i)(x - 8i) D) f(x) = x^2(8x + i)^2

OpenStudy (anonymous):

I think it is C. And I know, right @Loser66 ? I feel so dumb. I hate online school.

OpenStudy (mathstudent55):

Start by factoring out a common factor out of x^4 + 64x^2

OpenStudy (anonymous):

At first I thought it was b, but then i factored out that answer choice, and it didn't work. Maybe I screwed up...

OpenStudy (mathstudent55):

\(f(x) = x4 + 64x2 \) \(f(x) = x^2(x^2 + 64)\) \(f(x) = x^2(x^2 - (-64))\) \(f(x) = x^2(x + \sqrt{-64})(x - \sqrt{-64})\) \(f(x) = x^2(x + 8i)(x - 8i)\)

OpenStudy (anonymous):

Oh, I was right. :p But I didn't figure it out that way. How did you know to do the square root over the -64's, or how would you have if that hadn't been the only choice, seeing that 8i is in every answer choice.

OpenStudy (mathstudent55):

Now they are all linear factors: \(f(x) = x \cdot x(x + 8i)(x - 8i)\)

OpenStudy (mathstudent55):

If the last factor after factoring the x^2 term were x^2 - 64, that's the difference of 2 squares which factors into a the product of a sum and a difference. Instead, we had x^2 + 64. We can write an addition as a subtraction of the opposite, so x^2 + 64 = x^2 - (-64).

OpenStudy (anonymous):

Okay. I got this one. I have two more. Lol.

OpenStudy (mathstudent55):

Sorry, gtg.

OpenStudy (anonymous):

Okay.. :( God bless, and thanks for all your help.

OpenStudy (mathstudent55):

You're welcome.